2. Convert the following temperatures to Kelvin scale.
(a) 25°C
(b) 373°C

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) 25°C = (25 + 273) K = 298 K (b) 373°C = (373 + 273) K = 646 K